six years ago Jun was 4 times as old as Joan. In four years, he would be twice as old as Joan. How old are they now?
Answer by sudhanshu_kmr(1152)

Let present age of Jun and Joan are x and y respectively.
their ages six years ago: (x-6) and (y-6)
after four years : (x+4) and (y+4)
according to question
(x-6) = 4(y-6) ......(1)
and (x+4) = 2(y+4) ...(2)
using (1) we got, x-4y = -18
and using (2) we got x- 2y = 4

after solving these two equations...
y = 22/2 = 11

x = 4 + 2*11 = 26

so, Jun age is 26 years, Joan age is 11 years.